OpenStudy (anonymous):

may you check my answers? @jim_thompson5910

OpenStudy (anonymous):

i get that too

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[8^2\times 3.5^2=784\]

jimthompson5910 (jim_thompson5910):

each side is 3.5*8 = 28 so the total area is 28*28 = 784 you got it correct

OpenStudy (anonymous):

Thank you! :)